OpenStudy (anonymous):

Which of the following represents 9 x to the 3 tenths power in radical form?

OpenStudy (anonymous):

\[(9x)^{3 \over 10} = ((9x)^3)^{1 \over 10} = \sqrt[10]{(9x)^3}\]and if you want the inside "reduced", it's\[\sqrt[10]{729x^3}\]